# thanks: These authors contributed equallythanks: These authors contributed equally

### v2

@@ Abstract
Figure 2: (a,b) The evolution of the Liouvillean spectrum, $\mathrm{Im}(\lambda)$ and $\mathrm{Re}(\lambda^{\prime}/\gamma)$, as the dephasing rate $\gamma$ is varied. For improved visualization, the imaginary eigenvalues in panel (a) have been offset within a small window. Red and blue dots in (a) indicate the scarred eigenvalue spacings for the dephasing-free case. (c) Analytical prediction of LPTS breaking points $\gamma_{\star}^{\mathrm{S(1,2,3)}}$ (black stars) based on perturbation theory are in good agreement with the numerical data (see text for details). (d) The average derivative $\mathrm{Re}(\bar{v})$ as a function of the dephasing rate $\gamma$ for QMBS and thermal states. (e) The average spacing $\bar{\delta}$ between the scarred and thermal eigenenergies in the dephasing-free case, plotted as a function of the Hilbert space dimension $D$. All data is for the SCL model with couplings $J_{\mathrm{h}}\!=\!0.66$, $J_{\mathrm{x},j}\!=\!0.1$. Panels (a)-(d) are for system size $L\!=\!10$ with open boundary conditions.
LPTS breaking in perturbation theory—To understand the spectra in Fig. 2, we have developed a Liovillean degenerate perturbation theory valid for small $\gamma$, see SM [87] for details. The Hamiltonian of the scar subspace can be rewritten as $\hat{S}\!=\!\hat{S}_{0}+\hat{\Sigma}$, where $\hat{\Sigma}$ is the self-energy from the thermal subspace. If we neglect the coupling between the bath and the thermal subsystem, the Liouvillean of approximate scar is expressed as
$\mathcal{S}\approx\mathcal{S}_{0}+\mathcal{E}+\gamma_{\mathrm{eff}}\mathcal{D},\quad\quad\mathcal{E}={-i\big{(}\hat{\Sigma}\otimes\hat{I}-\hat{I}\otimes\hat{\Sigma}\big{)}},$ (4)
where the self-energy superoperator $\mathcal{E}$ is calculated using a non-equilibrium Green’s function and $\gamma_{\mathrm{eff}}$ is model-dependent [87].
For the SCL model, the scar states are simply the Dicke states $|E_{n_{\mathrm{S}}}\rangle$ of a free collective spin of magnitude $N/2$ [87]. The eigenmodes of the dephasing-free Liouvillean $\mathcal{S}_{0}$ are $|\lambda_{(l,s)}^{(0)}\rangle\!\rangle\!=\!|E_{l+s}\rangle\!\otimes\!|E_{s}\rangle$, corresponding to eigenvalues $\lambda_{(l,s)}^{(0)}\!=\!i(E_{l+s}\!-\!E_{s})\!=\!2il$ with indices $l\!=\!-N,1\!\!-\!N,\cdots,N$ and $s\!=\!0,1,\cdots,N\!\!-\!|l|$ labeling the eigenmodes of the scar subspace. For clarification, as illustrated in Fig. 1, the layer indices $l$ are divided into two classes: $l^{\prime\prime}$ for the sector with LPTS breaking and $l^{\prime}$ for the sector without LPTS breaking. In first-order perturbation theory, the corrections to the traceless Liouvillean eigenvalues for the layer $l$ are obtained by diagonalizing the projection of $\left(\mathcal{E}\!+\!\gamma_{\mathrm{eff}}\mathcal{D}^{\prime}\right)$ to the degenerate subspace of the $l$th layer. For the finite-size SCL model in Fig. 2, the perturbed eigenvalues can be expressed in closed form [87], and we plot them against the numerical result in Fig. 2(a) and (c).
@@
The analytical prediction of $\gamma_{\star}^{\mathrm{S}(1,2,3)}$ in perturbation theory aligns well with the numerical results [Figs. 2(a)-(c)], implying that LPTS breaking in QMBS Liouvillean spectra originates from the competition between dissipation $\gamma_{\mathrm{eff}}\mathcal{D}^{\prime}$ and the self-energy $\mathcal{E}$. Hence, we can estimate the LPTS breaking point for QMBS Liouvillean eigenvalues:
@@
which follows from the operator norm of $\mathcal{D}^{\prime}$ being extensive in system size, while $\|\hat{\Sigma}\|$ is independent of $L$ [87]. This estimate is consistent with Fig. 2(e), where we numerically extract the scaling behavior of the average spacing $\bar{\delta}$ for scar and thermal eigenvalues in the dephasing-free case ($\gamma\!=\!0$) as a function of the Hilbert space dimension $D$. The typical spacing for thermal eigenvalues scales as $\bar{\delta}\!\sim\!LD^{-2}$, while the scar $\bar{\delta}^{\mathrm{S}}$ is nearly independent of $D$. Then, the heuristic estimate of the typical LPTS breaking point, $\bar{\gamma}_{\star}^{\mathrm{S}}\!\sim\!\bar{\delta}^{\mathrm{S}}\|\mathcal{D}^{\prime}\|^{-1}$, is in agreement with Eq. (5). This is in stark contrast with the exponential scaling $\bar{\gamma}_{\star}\!\sim\!D^{-2}$ of thermal eigenvalues [82]. Thus, while $\bar{\gamma}_{\star}$ is prohibitively small for a generic many-body system, the presence of both QMBSs and $\mathcal{T}_{-}$ symmetry gives rise to LPTS breaking points that scale polynomially with system size and have robust experimental signatures, as shown below.
@@
Figure 3: The global fidelity density at the first revival, $\ln(F_{1})/L$, (a) and the decay coefficient $\beta$ of density imbalance (b) for the SCL model as a function of dephasing $\gamma$. The orange and blue shaded areas represent before and after the LPTS breaking point $\gamma_{\star}^{\mathrm{S}}$, respectively. For $\gamma\!<\!\gamma_{\star}^{\mathrm{S}}$, both the fidelity density and imbalance decay are well-converged in system size, suggesting the robustness of scar signatures over a finite range of $\gamma L$. By contrast, in the exact scar case with $J_{\mathrm{x},j}\!=\!0$, the imbalance decay is highly sensitive to $\gamma$ across the whole range. All data is for $J_{\mathrm{h}}\!=\!0.66$, $J_{\mathrm{x},j}\!\in\![0,0.2]$ drawn from a uniform distribution. Parameters $(\beta,\omega)$ are obtained by fitting the imbalance dynamics over the time period $[0,3.2]$. These results are obtained by TEBD algorithm based on ITensor library [89].
@@
Dynamical signatures—We demonstrate that weak dependence of $\bar{\gamma}_{\star}^{\mathrm{S}}$ on system size allows for QMBS signatures to persist in open-system dynamics. The time evolution is calculated from the Lindblad equation in the Markovian approximation [85] as $\mathrm{d}|\rho(t)\rangle\!\rangle/\mathrm{d}t\!=\!\mathcal{L}|\rho(t)\rangle\!\rangle$. We integrate this equation starting from the initial state $|\rho_{0}\rangle\!\rangle\!=\!|\rho_{\Pi}\rangle\!\rangle$ and characterize its subsequent time evolution respectively by computing the fidelity and the density imbalance,
$\begin{split}F(t)&=\big{|}\langle\!\langle\rho_{0}|\rho(t)\rangle\!\rangle\big{|}=\big{|}\sum_{k}\langle\!\langle\rho_{0}|\lambda_{k}\rangle\!\rangle^{2}e^{\lambda_{k}t}\big{|},\\ I(t)&=\frac{1}{L}\sum_{j}\big{\langle}\!\big{\langle}\hat{\sigma}^{z}_{j}\otimes\hat{\sigma}^{z}_{j}\big{\rangle}\!\big{\rangle}_{0}\big{\langle}\!\big{\langle}\hat{\sigma}^{z}_{j}\otimes\hat{\sigma}^{z}_{j}\big{\rangle}\!\big{\rangle}_{t},\end{split}$ (6)
where $\langle\!\langle\hat{\sigma}^{z}_{j}\!\otimes\!\hat{\sigma}^{z}_{j}\rangle\!\rangle_{t}\!=\!\langle\!\langle\rho(t)|\hat{\sigma}^{z}_{j}\!\otimes\!\hat{\sigma}^{z}_{j}|\rho(t)\rangle\!\rangle$. As the fidelity dynamics directly corresponds to the Liouvillean spectrum, the variance of fidelity density at the first revival point $\ln(F_{1})/L$ can manifest the Liouvillean spectrum transition. Furthermore, the imbalance dynamics takes the form of a damped oscillation, which can be fit using $\bar{I}(t)\!=\!\exp(-\beta t)\cos(\omega t)$ over short time periods. In Fig. 3, $\ln(F_{1})/L$ for the SCL model with dephasing is shown in panel (a), while the extracted imbalance parameters ($\omega$, $\beta$) are plotted in panel (b). Below the LPTS breaking point, both quantities are well-converged and weakly depend on $\gamma$. For exact scars ($J_{\mathrm{x},j}\!=\!0$) in Fig. 3(b), the decay exponent $\beta$ shows a power-law growth, while for approximate scar, $\beta$ is nearly constant before the breaking point. As the dephasing rate further increases, the behavior of approximate and exact scars become similar.
@@
Conclusions—By analyzing the Liouvillean spectra and dynamics in the presence of dephasing, we demonstrated that QMBSs can undergo a dissipative transition that resembles LPTS breaking. This behavior distinguishes approximate scars from thermal states; the former exhibit a certain degree of robustness to a dephasing dissipation, reminiscent of their behavior under closed system dynamics when prepared in a thermal Gibbs state [90]. We note that quantum noise is crucial for observing the QMBS LPTS breaking phenomena as the latter do not persist under a non-Hermitian Hamiltonian approximation [87].
@@
We supported our conclusions using the SCL model realized in superconducting platform [39]. In Appendix, we perform the same analysis for the PXP model of Rydberg atom arrays, finding a similar transition in its Liouvillean spectrum as a function of pure dephasing. While the PXP model lacks the $\mathcal{T}_{-}$ symmetry [Eq. (3)], it displays similar (albeit approximate) LPTS breaking and similar robustness of QMBS signatures. Notably, these two platforms allow for controlled dephasing noise [91, 92], indicating experimental implications for observing imbalance dynamics by tuning the dephasing rate. Further evidence of LPTS breaking for the 1D tilted Fermi-Hubbard model [35], and ladder models with Hilbert space fragmentation [93] is presented in the SM [87]. The last example suggests that the phenomenon of spontaneous LPTS breaking may be a general feature of systems that weakly break ergodicity, extending beyond QMBSs. Furthermore, it remains unclear whether the stable dynamics present in systems beyond those with QMBSs and how to engineer the stability of coherent quantum dynamics. These deserve more systematic investigations beyond the scope of the present work. Another open question concerns the effect of other types of dissipation on LPTS breaking and their underlying mechanisms. Finally, in closed systems, QMBSs have been proposed for quantum sensing [64], and it would be interesting to explore the sensitivity of such applications to dephasing noise.
Acknowledgements.
@@
Acknowledgments—We thank Profs. J.-S. Jin, Z.-W. Zhou, W.-G. Wang, X.-J. Liu, D.-L. Deng for helpful discussions. We also appreciate Dr. J. Ren discussion on TEBD simulation for quantum open systems. This work was supported by the Zhejiang Provincial Natural Science Foundation of China (No. LD25A050002), the National Natural Science Foundation of China (Nos. 12375021 and 12247101), the Fundamental Research Funds for the Central Universities (Grant No. lzujbky-2024-jdzx06), and the National Key Research and Development Program of China (No. 2022YFA1404203). Z.P. acknowledges support by the Leverhulme Trust Research Leadership Award RL-2019-015 and EPSRC Grant EP/Z533634/1. Statement of compliance with EPSRC policy framework on research data: This publication is theoretical work that does not require supporting research data. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP).
Appendix A Appendix: PXP Model
PXP model—A paradigmatic example of approximate scars is the PXP model that describes Rydberg atom arrays [14]. The PXP Hamiltonian is [94, 95]
$\hat{H}_{\mathrm{PXP}}=\Omega\sum^{L}_{j}\hat{P}_{j-1}\hat{\sigma}^{x}_{j}\hat{P}_{j+1},$ (7)
where $\hat{\sigma}_{j}^{x,y,z}$ are the Pauli matrices, $\hat{P}_{j}\!=\!(1-\hat{\sigma}^{z}_{j})/2$, and we set the Rabi frequency to $\Omega\!=\!1$. The model features revival dynamics when the system is quenched from the initial state $|\mathbb{Z}_{2}\rangle\!\equiv\!|{\uparrow}{\downarrow}{\uparrow}{\downarrow}{\cdots}\rangle$ [15, 96]. Below we demonstrate that the PXP model with pure dephasing exhibits similar behavior to the SCL model.
PXP spectral transition—Figures 4(a)-(c) illustrate the Liouvillean spectra for the PXP model for several dephasing rates. These results are qualitatively similar to the SCL model in Fig. 2: for weak dephasing $\gamma\!=\!0.0004$, the scar eigenvalues cluster around $\mathrm{Re}\big{(}\lambda^{\prime}\big{)}\!=\!0$, while at larger dephasing, $\gamma\!=\!0.04$, these eigenvalues pair-wisely shift to both sides along the real axis. The main difference is that the PXP model does not possess the $\mathcal{T}_{-}$ symmetry in Eq. (3). Hence, its Liouvillean spectrum, even at small $\gamma\!=\!0.0004$, does not have an exact $D_{2}$ symmetry and LPTS, strictly speaking, is already broken at that point. However, the clustering of the scar eigenvalues around the imaginary and real axes at small $\gamma$ suggests that their movement is largely similar to the SCL model. We next show that scar eigenvalues of the PXP Liouvillean indeed undergo sharp jumps upon sweeping the dephasing rate, similar to the SCL model.
Figure 4: (a-c) The Liouvillean eigenspectra of the PXP model for dephasing strengths $\gamma\!=\!0$, $0.0004$, and $0.04$. Red and blue dots represent scar (S) eigenvalues, while yellow and gray ones represent thermal (T) eigenvalues. These are further classified by momenta $q\!=\!0$ and $\pi$ under translations. The inset in (a) shows the average scar eigenvalue spacing, $\bar{\delta}$, at zero dephasing, plotted as a function of the corresponding sector dimension $D$. The right panels in (b)-(c) show the overlaps of Liouvillean eigenmodes with $|\rho_{\mathrm{Z_{2}}}\rangle\!\rangle$, with black circles highlighting local maxima. (d)-(e) The evolution of the Liouvillean spectrum, $\mathrm{Im}(\lambda)$ and $\mathrm{Re}(\lambda^{\prime}/\gamma)$, as the dephasing rate $\gamma$ is varied. The color intensity signifies the extent of real-axis difference. For improved visualization, the imaginary values in panel (d) have been shifted within a small window. Red dots indicate the scarred eigenvalue spacings for the dephasing-free case. (f) The average derivative $\mathrm{Re}(\bar{v})$ as a function of the dephasing rate $\gamma$ for the QMBSs and the thermal eigenvalues. All data is obtained by exact diagonalization of the PXP model for a system size of $L\!=\!10$ and with periodic boundary conditions.
Figures 4(d) and (e) illustrate the imaginary part of eigenvalues and real part of normalized eigenvalues $\lambda^{\prime}/\gamma$ as a function of $\gamma$. Several eigenvalues with momentum $q=0$ remain stationary at weak dephasing, while some pairs of eigenvalues abruptly bifurcate into two branches at specific dephasing rates $\lambda_{\star}^{(1,2,3)}$. In contrast, no such LPTS breakings occur in the momentum sector $q=\pi$. The LPTS breaking can be seen in the average movement velocity $\mathrm{Re}(v)$ as a function of $\gamma$ in Fig. 4(f). For the $q=0$ momentum sector of the PXP model, a few distinct peaks align with the fast breaking velocities in Figs. 4(d)-(e) and the breaking points $\gamma_{\star}^{(1,2,3)}$ correspond to the front of three peaks. There are no such abrupt peaks in $q=\pi$ sector of QMBSs. Moreover, the average movement velocity of thermal eigenvalues become more stable with increasing $\gamma$, implying no critical points.
The typical LPTS breaking point $\bar{\gamma}_{\star}^{\mathrm{S}}$ for PXP scar eigenmodes can be estimated from the spacings $\bar{\delta}$ of nondegenerate scar eigenvalues at $\gamma\!=\!0$ according to the same expression, $\bar{\gamma}_{\star}^{\mathrm{S}}\!\sim\!\bar{\delta}\;\|\mathcal{D}^{\prime}\|^{-1}$, used previously for the SCL model in the main text. As confirmed in the inset of Fig. 4(a), $\bar{\delta}\!\sim\!LD^{-2}$ for thermal eigenvalues, while the scar $\bar{\delta}$ is nearly independent of $D$. Thus, the typical critical point for PXP scar eigenmodes also scales as $\bar{\gamma}_{\star}^{\mathrm{S}}\!\sim\!L^{-1}$, consistent with the approximate scars in the SCL model.
Figure 5: The global fidelity density at the first revival, $\ln(F_{1})/L$ (a) and the decay coefficient $\beta$ of density imbalance (b) for the PXP model as a function of dephasing $\gamma$. The orange and blue areas represent before and after the LPTS breaking point $\gamma_{\star}^{\mathrm{S}}$, respectively. Parameters $(\beta,\omega)$ are obtained by fitting the imbalance dynamics over the time period $[0,5]$. These results are obtained by exact diagonalization.
PXP dynamical signatures—Similar to the SCL model in the main text, we check for the robustness of scar signatures in the dynamics by performing the time evolution generated by the Lindblad equation for the PXP model with pure dephasing. Here we take as initial state $|\rho_{\mathrm{Z}_{2}}\rangle\!\rangle\!=\!|\mathbb{Z}_{2}\rangle\!\otimes\!|\mathbb{Z}_{2}\rangle$ and the resulting fidelity density at the first revival, $\ln F_{1}/L$, as well as the imbalance decay $\beta$ and frequency $\omega$, are shown in Fig. 5. These results mirror those for the SCL model, in particular $\beta$ has a similarly weak dependence on $\gamma$ as long as $\gamma\!<\!\gamma_{\star}^{\mathrm{S}}$.
To confirm that these results are due to approximate scars, we consider a deformation of the PXP model [17, 19, 63] for which QMBSs assume a nearly exact form:
$\hat{H}_{\mathrm{DPXP}}=\hat{H}_{\mathrm{PXP}}+\sum^{L}_{j}\sum^{L/2}_{d=2}J_{\mathrm{d}}\hat{P}_{j-1}\hat{\sigma}^{x}_{j}\hat{P}_{j+1}\big{(}\hat{\sigma}^{z}_{j-d}+\hat{\sigma}^{z}_{j+d}\big{)}.$ (8)
At the optimal coupling strength $J_{\mathrm{d}}\!=\!0.051(\phi^{d-1}\!+\!\phi^{1-d})^{-2}$, with $\phi=(1+\sqrt{5})/2$, the scar subspace contains $L+1$ eigenstates that are almost perfectly decoupled from the thermal bulk of the spectrum [17]. This DPXP model indeed behaves similarly to the exact scars in the SCL model with $J_{\mathrm{x},j}\!=\!0$, e.g., the imbalance decay coefficient $\beta$ exhibits a similar power-law dependence on $\gamma$ in both cases. The similarity extends to the Liouvillean spectra, as the DPXP model also exhibits LPTS breaking for infinitesimal $\gamma$, see SM [87].
@@ Appendix A Appendix: PXP Model
“Liouvillean Spectral Transition in Noisy Quantum Many-Body Scars”
@@
Appendix S1 A. Liouvillean perturbation theory

### v3

@@ Abstract
Figure 2: (a) The evolution of the Liouvillian spectrum in the CL model, Eq. (Liouvillian Spectral Transition in Noisy Quantum Many-Body Scars), with couplings $J_{\mathrm{h}}\!=\!0.66$, $J_{\mathrm{x},j}\!=\!0.1$ and system size L=10L\!=\!10 with open boundary conditions. We plot $\mathrm{Im}(\lambda)$ and $\mathrm{Re}(\lambda^{\prime}/\gamma)$, as the dephasing rate $\gamma$ is varied. For improved visualization, the imaginary eigenvalues have been offset within a small window. Analytical prediction of perturbation theory (lines) are in good agreement with the numerical data, in particular for the spectral transition points $\gamma_{\star}^{\mathrm{S(1,2,3)}}$ (black stars). (b) The average spectral transition point, $\bar{\gamma}_{\star}^{S}\,L$, as a function of the system size $L$ obtained in perturbation theory. (c) The average spacing $\bar{\delta}$ between the scarred and thermal eigenenergies in the dephasing-free case, plotted as a function of the Hilbert space dimension $D$.
To understand the spectral evolution in Fig. 2(a), we have developed a Liouvillian degenerate perturbation theory valid for small $\gamma$. If we neglect the coupling between the bath and the thermal subsystem, the effective Liouvillian for the scar subspace is $\mathcal{L}_{\mathcal{S}}\approx\mathcal{S}_{0}+\mathcal{E}+\gamma_{\mathrm{eff}}\mathcal{D}^{\prime}$, where $\mathcal{E}={-i\big(\hat{\Sigma}\otimes\hat{I}-\hat{I}\otimes\hat{\Sigma}\big)}$ is the self-energy superoperator that includes the effect of thermal bulk [86]. The effective dephasing $\gamma_{\mathrm{eff}}=2\gamma$ accounts for the magnitude of collective spin being half the system size, $N=L/2$.
In the regime $\|\mathcal{E}+\gamma_{\mathrm{eff}}\mathcal{D}^{\prime}\|\ll\|\mathcal{S}_{0}\|$, the traceless Liouvillian eigenvalues for the $l$th layer are modified with respect to Eq. (4) into $\lambda^{\prime}_{(l,s)}\approx\lambda^{(0)}_{(l,s)}+\lambda^{(1)}_{(l,s)}$, where the first-order corrections $\lambda^{(1)}_{(l,s)}$ are obtained by diagonalizing the projection of $\left(\mathcal{E}\!+\!\gamma_{\mathrm{eff}}\mathcal{D}^{\prime}\right)$ to the degenerate subspace of the $l$th layer. The projected matrix consists of an imaginary diagonal matrix $\mathcal{E}_{l}$ and a real shift matrix $\mathcal{D}_{l}^{\prime}$:
$\mathcal{E}_{l}$ $=i\times\mathrm{Diag}\Big[\Sigma_{(l,0)},\ \Sigma_{(l,1)},\ \cdots,\Sigma_{(l,N-|l|)}\Big],$ (5)
$\mathcal{D}_{l}^{\prime}$ $=\sum_{s=0}^{N-|l|-1}d_{(l,s)}\Big[\ \big|\lambda_{(l,s)}^{(0)}\big\rangle\!\big\rangle\big\langle\!\big\langle\lambda_{(l,s+1)}^{(0)}\big|+\mathrm{h.c.}\Big],$
with $\Sigma_{(l,s)}=\big\langle E_{s}^{\mathrm{S}}\big|\hat{\Sigma}\big|E_{s}^{\mathrm{S}}\big\rangle-\big\langle E_{s+l}^{\mathrm{S}}\big|\hat{\Sigma}\big|E_{s+l}^{\mathrm{S}}\big\rangle$ and $d_{(l,s)}=\sqrt{\left(s\!+\!1\right)\left(N\!-\!s\right)\left(|l|\!+\!s\!+\!1\right)\left(N\!-\!|l|\!-\!s\right)}/N$ [86]. By diagonalizing $\left(\mathcal{E}_{l}\!+\!\gamma_{\mathrm{eff}}\mathcal{D}_{l}^{\prime}\right)$, we obtain the perturbed eigenvalues for the $l$th layer and plot them against the numerical results for the same parameters in Fig. 2(a).
@@
The perturbation theory results in Fig. 2(a) compare well with the numerics for both the imaginary and the real parts of the Liouvillian spectrum. The analytical prediction of $\gamma_{\star}^{\mathrm{S}(1,2,3)}$ implies that the spectral transition in QMBS Liouvillian spectra originates from the competition between dissipation $\gamma_{\mathrm{eff}}\mathcal{D}^{\prime}$ and self-energy $\mathcal{E}$. While different QMBS eigenvalues undergo bifurcations at generally different values of $\gamma_{\star}^{\mathrm{S}(1,2\ldots)}$, the mean value $\bar{\gamma}_{\star}^{\mathrm{S}}$ has a robust universal scaling with $L$ and determines the dynamical signatures studied below. Hence, we estimate the spectral transition point for QMBS Liouvillian eigenvalues:
@@
which follows from the operator norm of $\mathcal{D}^{\prime}$ being extensive in system size, while $\|\mathcal{E}\|$ is independent of $L$ [86]. This estimate, shown in Fig. 2(b), is consistent with Fig. 2(c), where we numerically extract the average spacing $\bar{\delta}$ for scar and thermal eigenvalues in the dephasing-free case ($\gamma\!=\!0$) as a function of the Hilbert space dimension $D$. The typical spacing for thermal eigenvalues scales as $\bar{\delta}\!\sim\!LD^{-2}$, while the scar $\bar{\delta}^{\mathrm{S}}$ is nearly independent of $D$. Then, the heuristic estimate of the typical spectral transition point, $\bar{\gamma}_{\star}^{\mathrm{S}}\!\sim\!\bar{\delta}^{\mathrm{S}}\|\mathcal{D}^{\prime}\|^{-1}$, is in agreement with Eq. (6). This is in stark contrast with the exponential scaling $\bar{\gamma}_{\star}\!\sim\!D^{-2}$ of thermal eigenvalues [82]. Thus, the presence of QMBSs gives rise to spectral transition points that scale polynomially with system size and have robust dynamical signatures, as shown next.
@@
Figure 3: The time evolution of the fidelity $F(t)$ (a) and the density imbalance $I(t)$ (b) for two dephasing rates at system size $L=16$. The black circles in Fig. 3 (a) represent the first peak of $F(t)$. The black dash-dotted lines in Fig. 3(b) represent the fit of $\bar{I}(t)$ (see text). The fidelity density at the first revival, $\ln(F_{1})/L$ (c) and the decay coefficient $\beta$ of density imbalance (d) for the CL model as a function of dephasing $\gamma$. For $\gamma\!<\!\gamma_{\star}^{\mathrm{S}}$, both the fidelity density and imbalance decay are well-converged in system size, suggesting the robustness of scar signatures over a finite range of $\gamma L$. By contrast, in the exact scar case with $J_{\mathrm{x},j}\!=\!0$, the imbalance decay is highly sensitive to $\gamma$ across the whole range. All data is for $J_{\mathrm{h}}\!=\!0.66$, $J_{\mathrm{x},j}\!\in\![0,0.2]$ drawn from a uniform distribution. Parameters $(\beta,\omega)$ are obtained by fitting the imbalance dynamics over the time period $[0,3.2]$. These results are obtained by TEBD algorithm based on ITensor library [88].
@@
Dynamical signatures—We demonstrate that weak dependence of $\bar{\gamma}_{\star}^{\mathrm{S}}$ on system size allows for QMBS signatures to persist in open-system dynamics defined by the evolution equation $\mathrm{d}|\rho(t)\rangle\!\rangle/\mathrm{d}t\!=\!\mathcal{L}|\rho(t)\rangle\!\rangle$ [87]. We integrate this equation starting from the initial state $|\rho_{0}\rangle\!\rangle\!=\!|\rho_{\Pi}\rangle\!\rangle$ and characterize subsequent time evolution by computing the fidelity and the density imbalance,
$F(t)=\Big|\sum_{k}\langle\!\langle\rho_{0}|\lambda_{k}\rangle\!\rangle^{2}e^{\lambda_{k}t}\Big|,\quad I(t)=\frac{1}{L}\sum_{j}\langle\!\langle\mathcal{Z}_{j}\rangle\!\rangle_{0}\langle\!\langle\mathcal{Z}_{j}\rangle\!\rangle_{t},$ (7)
where $\langle\!\langle\mathcal{Z}_{j}\rangle\!\rangle_{t}=\!\langle\!\langle\rho(t)|\hat{\sigma}^{z}_{j}\!\otimes\!\hat{\sigma}^{z}_{j}|\rho(t)\rangle\!\rangle$. As the fidelity dynamics directly corresponds to the Liouvillian spectrum, the variance of fidelity density at the first revival point $\ln(F_{1})/L$, illustrated in Fig. 3(a), can reveal the Liouvillian spectrum transition. Furthermore, the imbalance dynamics takes the form of a damped oscillation [Fig. 3(b)], which we fit using $\bar{I}(t)\!=\!\exp(-\beta t)\cos(\omega t)$ to extract the decay rate $\beta$ and revival frequency $\omega$. In practice, we found it sufficient to fit the data up to the first revival period [86].
In Fig. 3(c), we plot $\ln(F_{1})/L$ as a function of $\gamma L$, while the extracted imbalance parameters ($\omega$, $\beta$) are plotted in Fig. 3(d). Below the spectral transition point, both quantities are well-converged and weakly depend on $\gamma L$, which is indeed consistent with the spectral transition results in Fig. 1 and perturbation theory estimate in Eq. (6).
The dissipative dynamics in Fig. 3 highlight the difference between the QMBS systems where the su(2) spectrum generating algebra is exact ($J_{\mathrm{x},j}\!=\!0$) versus only approximately obeyed ($J_{\mathrm{x},j}\!\neq\!0$) within the scar subspace. In Fig. 3(c)-(d), we notice that the fidelity density and decay exponent $\beta$ in the exact scar case both display power-law dependence on $\gamma L$, in stark contrast with the plateau observed for $J_{\mathrm{x},j}\!\neq\!0$. Thus, only approximate QMBSs undergo the spectral transition at a non-zero $\gamma_{\star}^{\mathrm{S}}$. This difference in behavior is due to the exact energy spacing at $J_{\mathrm{x},j}\!=\!0$, which leads to degeneracies on the imaginary axis in Eq. (4). Hence, the Liouvillian eigenvalues of exact scars jump into the complex plane for any $\gamma\!\neq\!0$, as confirmed in the SM [86]. However, as the dephasing rate in Fig. 3(c)-(d) increases beyond $\gamma_{\star}^{\mathrm{S}}$, the behaviors of approximate and exact scars become similar.
The nature of the spectral transition.—In generic chaotic models (without QMBSs), the LPTS breaking can be understood from the symmetry classification developed in Ref. [89]. In particular, the existence of a superoperator $\mathcal{T}_{-}$ that obeys $\mathcal{T}_{-}\mathcal{L}\mathcal{T}_{-}^{-1}=$-$\mathcal{L}$ is responsible for the dihedral symmetry of the Liouvillian spectrum and $\mathbb{P}\mathbb{T}$ breaking. At $\gamma\!=\!0$, $|-(\lambda^{\prime})^{\ast}\rangle\!\rangle$ and $|\lambda^{\prime}\rangle\!\rangle$ represent the same state which lies on the imaginary axis. Beyond a certain $\gamma$, $\mathbb{P}\mathbb{T}|\lambda^{\prime}\rangle\!\rangle\!=\!|\!-\!(\lambda^{\prime})^{\ast}\rangle\!\rangle$ continues to hold, but $\lambda^{\prime}$ and $(\lambda^{\prime})^{\ast}$ are no longer the same. Thus, after this critical point, pairs of degenerate eigenvalues move symmetrically away from the imaginary axis along the real axis, resulting in LPTS breaking [82].
One may wonder if a similar, symmetry-based argument can be applied to the QMBS spectral transition observed in Fig. 1 above. For the CL model, there indeed exists an analogous operator $\mathcal{T}_{-}$ (see [86] for its explicit form). However, the spectral transition in the CL model only occurs in inversion symmetry sector p=1(+1)p\!=\!-1(+1), corresponding to odd (even) $N$. See N=5N\!=\!5 case in Fig. 1. In contrast, the symmetry is already broken for infinitesimal $\gamma$ in p=+1(1)p\!=\!+1(-1) sector, also correlating with even (odd) $N$. Thus, $\mathcal{T}_{-}$ by itself does not fully account for the observations in Fig. 1. As a more transparent example, in End Matter, we perform the same analysis for the PXP model of Rydberg atom arrays, finding a similar transition in its Liouvillian spectrum as a function of dephasing. The PXP model manifestly lacks the $\mathcal{T}_{-}$ symmetry [86], while it displays a spectral transition similar to Fig. 1 and robustness of QMBS dynamical signatures similar to Fig. 3. Thus, we conclude the QMBS spectral transition is not due to symmetry breaking, but due to the smallness of matrix elements connecting the QMBS subspace with thermal bulk of the spectrum.
@@
Conclusions—By analyzing the Liouvillian spectra and dynamics in the presence of dephasing, we demonstrated that QMBSs can undergo a dissipative transition that resembles LPTS breaking. This behavior distinguishes approximate scar states from thermal states; the former exhibit a certain degree of robustness to a dephasing dissipation, reminiscent of their behavior under closed system dynamics when prepared in a thermal Gibbs state [90]. We note that quantum noise is crucial for observing the QMBS Liouvillian spectral transition phenomena as the latter do not persist under a non-Hermitian Hamiltonian approximation [86].
@@
We supported our conclusions using the CL and PXP models realized in superconducting and Rydberg platforms. Notably, these two platforms allow for controlled dephasing noise [91, 92] and can be used test our predictions for the imbalance dynamics by tuning the dephasing rate. Further numerical evidence for the Liouvillian spectral transition in the 1D tilted Fermi-Hubbard model [35], and ladder models with Hilbert space fragmentation [93] is presented in the SM [86]. All of these models are described by the similar effective Hamiltonian $\hat{H}=\hat{S}_{0}\oplus\hat{T}+\hat{V}$, suggesting broader applicability of our conclusions to models with this structure. Furthermore, the example of Hilbert space fragmentation suggests that the Liouvillian spectral transition may extend beyond QMBSs to other types of weak ergodicity breaking, warranting further study. Another open question concerns other types of dissipation beyond pure dephasing and their impact on the spectral transition. Finally, in closed systems, QMBSs have been proposed for quantum sensing [60], and it would be interesting to explore the sensitivity of such applications to dephasing noise.
@@
Acknowledgments—We thank Profs. Jiasen Jin, Zheng-Wei Zhou, Wen-Ge Wang, Xiong-Jun Liu, Dong-Lin Deng for helpful discussions. We also appreciate Dr. Jie Ren discussion on TEBD simulation for quantum open systems. This work was supported by the National Natural Science Foundation of China (Nos. 12375021 and 12247101), the Zhejiang Provincial Natural Science Foundation of China (No. LD25A050002), the National Key Research and Development Program of China (No. 2022YFA1404203) and the Fundamental Research Funds for the Central Universities (Grant No. lzujbky-2024-jdzx06). Z.P. acknowledges support by the Leverhulme Trust Research Leadership Award RL-2019-015 and EPSRC Grant EP/Z533634/1. Statement of compliance with EPSRC policy framework on research data: This publication is theoretical work that does not require supporting research data. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP).
@@
“Liouvillian Spectral Transition in Noisy Quantum Many-Body Scars”
@@
S1 A. Liouvillian perturbation theory